OpenStudy (aravindg):

prove : [tan theta/(1-cot theta)] +[cot theta/(1-tan theta)]=(1+tan theta +cot theta)

OpenStudy (anonymous):

(tan[1-tan]+cot[1-cot])/({1-cot}[1-tan])

OpenStudy (anonymous):

(tan-tan^2)+(cot-cot^2) --------------------- 2-cot-tan (cot*tan=1)

OpenStudy (aravindg):

??

OpenStudy (anonymous):

here i have taken x instead of theta for simplicity LHS = cot(x) / (1-tan(x)) + tan(x) / (1 - cot(x)) = (1/tan(x)) / (1 - tan(x)) + tan(x) / (1 - 1/tan(x)) = 1 / (tan(x)(1-tan(x)) + tan²(x) / (tan(x) - 1) = 1 / (tan(x)(1-tan(x)) - tan²(x) / (1- tan(x)) = (1 - tan³(x)) / (tan(x)(1-tan(x)) = (1 - tan(x))(1 + tan(x) + tan²(x)) / (tan(x)(1-tan(x)) = (1 + tan(x) + tan²(x) / tan(x) = cot(x) + 1 + tan(x) = 1 + tan(x) + cot(x) = RHS